# RigAnything: Template-Free Autoregressive Rigging for Diverse 3D Assets

### v1

@@ Abstract.
We present RigAnything, a novel autoregressive transformer-based model, which makes 3D assets rig-ready by probabilistically generating joints, skeleton topologies, and assigning skinning weights in a template-free manner. Unlike most existing auto-rigging methods, which rely on predefined skeleton template and are limited to specific categories like humanoid, RigAnything approaches the rigging problem in an autoregressive manner, iteratively predicting the next joint based on the global input shape and the previous prediction. While autoregressive models are typically used to generate sequential data, RigAnything extends their application to effectively learn and represent skeletons, which are inherently tree structures. To achieve this, we organize the joints in a breadth-first search (BFS) order, enabling the skeleton to be defined as a sequence of 3D locations and the parent index. Furthermore, our model improves the accuracy of position prediction by leveraging diffusion modeling, ensuring precise and consistent placement of joints within the hierarchy. This formulation allows the autoregressive model to efficiently capture both spatial and hierarchical relationships within the skeleton. Trained end-to-end on both RigNet and Objaverse datasets, RigAnything demonstrates state-of-the-art performance across diverse object types, including humanoids, quadrupeds, marine creatures, insects, and many more, surpassing prior methods in quality, robustness, generalizability, and efficiency. Please check our website for more details: https://www.liuisabella.com/RigAnything.
@@
conference: ; Feb; 2025ccs: Computing methodologies Animationccs: Computing methodologies Neural networks
@@ 1. Introduction
Figure 2. Skeleton generation on shapes obtained from real images, demonstrating that our method generalizes well to real data. Fig. 7 presents additional real results.
@@
Specifically, we represent the tree-structured skeleton as a sequence by ordering the joints in a breadth-first search (BFS) order, where each joint is defined by a 3D position and a parent index. This autoregressive formulation is particularly suited for skeleton prediction, as it addresses the inherent ambiguity in joint configurations by representing them as a probabilistic distribution. Additionally, by sequentially generating joints and connections without relying on a predefined template, the model supports arbitrary skeleton structures and varying numbers of joints, enabling broad generalization across diverse object categories. Furthermore, while autoregressive models are traditionally designed to handle discrete values (Waswani et al., 2017; Brown et al., 2020; Radford et al., 2019), inspired by recent work utilizing autoregressive models for image generation (Li et al., 2024), we adopt a diffusion sampling process to predict the continuously valued joint positions, resulting in superior accuracy. Given the predicted skeleton, we infer the skinning weights by a pair-wise computation. We employ transformer blocks throughout the model to comprehensively capture the global shape structure, as well as the interdependence among all joints and their associated surface points.
Table 1. Feature comparison with other Auto-Rigging tools.
@@ 2.1. Automatic Rigging
Rigging is a fundamental technique for animation in computer graphics. Traditional automatic rigging methods, such as Pinocchio (Baran and Popović, 2007), rely on predefined skeletons and optimize their variations to fit a range of characters, with skinning weights determined by analyzing vertex-bone deformation relationships. However, the optimization process is computationally expensive and diminishes the generalizability. Recent advances in deep learning have improved the quality and adaptability of rigging. TARig (Ma and Zhang, 2023) utilizes a template with adaptive joints and a boneflow field to generate skeletons and skinning weights for humanoid characters. Li’s et al. (Li et al., 2021) leverages a predefined skeleton template for characters to learn rigging and proposes neural blend shape to enhance deformation quality. However, these methods are confined to humanoid characters in standard poses and rely heavily on predefined templates, limiting their robustness and generalization to diverse objects, poses, and skeleton topologies.
Figure 3. Pipeline of our method: The input shape and the previously predicted skeleton sequence are tokenized using two separate tokenizers. These tokens are processed through a chain of autoregressive transformer blocks with a hybrid attention mask. Shape tokens perform self-attention to capture global geometric information, while skeleton tokens attend to all shape tokens and use causal attention within themselves to maintain the autoregressive generation process. After the transformer blocks, a skinning module decodes shape tokens into skinning weights, a joint diffusion module samples the next joint position, and a connectivity module predicts the next joint’s connection to its preceding joint conditioned on the sampled next joint position from joint diffusion module.
@@ 2.2. Autoregressive Models for 3D
Autoregressive models are a powerful class of probabilistic models widely applied across domains such as natural language processing (Brown et al., 2020; Achiam et al., 2023; Radford et al., 2019) and computer vision (Esser et al., 2021; Parmar et al., 2018; Chen et al., 2020; Li et al., 2024). In 3D tasks, autoregressive models have also demonstrated remarkable potential in areas like shape generation (Yan et al., 2022; Mittal et al., 2022; Ibing et al., 2023; Cheng et al., 2022; Qian et al., 2024) and motion generation (Zhang et al., 2023; Rempe et al., 2021; Han et al., 2024). In 3D shape generation, methods mostly focus on designing effective representations for autoregressive modeling. ShapeFormer (Yan et al., 2022) introduces a sparse representation that quantizes non-empty voxel grids in a predefined order. AutoSDF (Mittal et al., 2022) takes a different approach by modeling the entire space and using randomized sampling orders to enable non-sequential modeling. Octree Transformer (Ibing et al., 2023) introduce octree-based hierarchical shape representations with adaptive compression, significantly reducing sequence lengths. Cheng et al. (Cheng et al., 2022) decompose point clouds into semantically aligned sequences. Argus3D (Qian et al., 2024) utilizes discrete representation learning on a latent vector and scales up the model to improve the quality and versatility of 3D generation. Similarly, autoregressive models have advanced 3D motion generation. T2M-GPT (Zhang et al., 2023) uses motion VQ-VAE and textural descriptions for human motion generation. HuMoR (Rempe et al., 2021) proposes hierarchical latent variables for realistic motion synthesis. AMD (Han et al., 2024) presents an autoregressive model that iteratively generates complex 3D human motions from long textual descriptions. In this paper, we pioneer the application of autoregressive models to the task of automatic rigging, marking a significant advancement in this domain.
@@ 3. Method
Our goal is to transform a given 3D shape into an animatable asset by generating a plausible skeleton and predicting the corresponding skinning weights. These enable the 3D asset to be articulated under joint transformations using Linear Blend Skinning (LBS). In this section, we first introduce our novel autoregressive model for skeleton generation in Sec. 3.1, followed by our approach to skinning weight prediction in Sec. 3.2, and conclude with a detailed description of the model architecture in Sec. 3.3.
@@ 3.1. Autoregressive Skeleton Prediction
Autoregressive Modeling. The key component of our method is an autoregressive model for the skeleton prediction to address the ambiguity of skeleton structures and eliminates the need for predefined templates. To convert the tree-structured skeleton to a sequence that can be effectively processed by the autoregressive model, we adopt the breadth-first search order (BFS) to serialize the skeleton to a list:
@@
where $j_{k}\in\mathbb{R}^{3}$ and $p_{k}\in\{1,...,K\}$ denote the 3D position and the parent index of the $k$-th joint respectively. As we adopt the BFS order, $p_{k}<k$ and the first element $(j_{1},p_{1})$ always represents the root joint. The order of joints at the same BFS depth level is non-deterministic. To resolve this ambiguity, we randomly sample the order during training and uses generative modeling to cover the uncertainty.
@@
Instead of directly modeling in the original joint space, we map all previously predicted joints and their corresponding parents into a higher-dimensional token space to enhance the model’s expressive capacity. This token space effectively represents the evolving state of the skeleton, capturing its structural and hierarchical information as new joints and connections are incrementally added. Similarly, a sequence of shape tokens is extracted to encapsulate the global structure of the input shape, providing consistent contextual information throughout the modeling process. Denoting the skeleton tokens as $T_{1:k-1}\in\mathbb{R}^{(k-1)\times d}$ and the shape token as $H\in\mathbb{R}^{L\times d}$, the prediction targets are reformulated as:
@@
We train a noise estimator $\epsilon_{\theta}$, conditioned on the diffusion time step $m$ and the context $Z$$\in\mathbb{R}^{\left(L+k-1\right)\times d}$, where
@@
capturing both the evolving skeleton state and the input shape. Here, Tr refers to the transformer blocks that are detailed in Sec. 3.3. The model takes the noisy joint $j^{m}$ as input and predicts the added noise $\epsilon$. The training objective is defined as:
@@ 3.3. Autoregressive Transformer Architecture
Our autoregressive modeling is anchored on a transformer-based architecture, which outputs the shape tokens $H\in\mathbb{R}^{L\times d}$ and skeleton tokens $T_{1:k}\in\mathbb{R}^{k\times d}$ ($0<k\leq K$) that serve as conditional inputs for the autoregressive modeling for skeleton prediction (Sec. 3.1) and skinning prediction (Sec. 3.2). The extraction of these tokens involve two steps: first, referred as the “tokenization” step, an initial shape token and skeleton tokens are lifted from the raw input, this step produces a higher dimensional vector that has sufficient capacity in preparation to capture richer information in the further processing steps in the transformer; subsequently, the transformer process the these tokens through a series of attention blocks with carefully crafted attention masking to obtain the final shape and skeleton tokens, which are finally used as inputs to the skeleton and skinning prediction modules in Sec. 3.1 and Sec. 3.2.
@@ 4.1. Implementation Details
The implemented transformer consists of 12 layers with a hidden dimension of 1024. Following the implementation in (Zhang et al., 2025), each transformer block incorporates a multi-head self-attention layer with 16 heads and a two-layered MLP with a hidden dimension of 4096 and a GeLU activation. We employ Pre-Layer Normalization, Layer Normalization (LN), and residual connections consistent with the reference implementation. During training, we employ a hybrid attention masking strategy: shape tokens perform self-attention to effectively capture geometric information, while skeleton tokens use causal attention, attending only to their ancestor skeleton tokens within the sequence to facilitate auto-regressive generation. Additionally, skeleton tokens attend to all shape tokens. During inference, the network processes shape tokens as input and generates skeleton tokens in an auto-regressive manner.
The joint diffusion process follows (Nichol and Dhariwal, 2021; Li et al., 2024), which has a cosine noise scheduler with 1000 training steps and 300 resampling steps during inference. The denoising MLP is conditioned on the transformer-outputted joint tokens, where these tokens are incorporated into the noise scheduler’s time embedding through AdaLN (Peebles and Xie, 2023) within the Layer Normalization layers.
The fusing module is a two-layer MLP with an input size of 3072 and hidden dimensions of 2048 and 1024. During inference, after obtaining the next joint position via diffusion sampling, a shape tokenizer generates a latent shape token (dimension 1024), which is concatenated with previous context tokens (dimension 1024) and positional embeddings. The fusing module’s output serves as the updated context for connectivity and skinning prediction. Both the connectivity and skinning modules share a similar architecture with the fusing module, except their input size is 2048.
@@ 4.2. Dataset
We utilize both the RigNet dataset (Xu et al., 2020) and the Objaverse dataset (Deitke et al., 2023). The RigNet dataset contains 2,354 high-quality 3D models with ground-truth skeleton and skinning annotations. The Objaverse dataset offers a large collection of 3D models with varying rigging quality. To ensure data reliability, we filtered out 9,686 models with consistent and accurate skeleton and skinning information. Our dataset spans a diverse range of categories, including bipedal, quadrupedal, avian, marine, insectoid, and manipulable rigid objects. For each model, we sample point clouds and face normals from the mesh surface.
@@
During training, we perform online augmentation to the input data by randomly deforming the input point clouds using the ground-truth skeleton and skinning. As shown in our ablation study in Sec. 4.4 and Fig. 6, this augmentation strengthens our method’s ability to generalize to objects in different poses.
4.3. Evaluation and Baseline Comparisons
4.3.1. Skeleton Prediction
We provide qualitative visualizations of the reconstructed skeletons in comparison with the ground truth and RigNet in Fig. 5. Our method demonstrates superior performance, producing more accurate and satisfying skeletons across various shape categories. In contrast, RigNet (Xu et al., 2020) struggles to recover reasonable skeletons for less common shapes, such as characters with tails or wings, and frequently generates an excessive number of joints. In comparison, our method generates a reasonable number of joints, with the reconstructed skeletons well-aligned to the underlying shape, ensuring better structural consistency and fidelity.
To quantitatively evaluate the performance of skeleton prediction, we measure the similarity between the predicted skeletons and the ground truth on the RigNet dataset using multiple metrics: Intersection over Union (IoU), Precision, and Recall for bone matching, as well as Chamfer distances for joints (CD-J2J), bone line segments (CD-B2B), and joint-to-bone line segments (CD-J2B). Table 2 presents a comparison with Pinocchio (Baran and Popović, 2007) and RigNet (Xu et al., 2020) across these metrics. The results show that our method significantly outperforms the baselines, producing skeletons that align more closely with the ground truth.
Table 2. Quantitative comparison of skeleton prediction on the RigNet dataset. Our predicted skeletons alignes better with the ground truth.
Pinocchio 0.365 0.387 0.359 0.072 0.055 0.047
RigNet 0.616 0.676 0.589 0.039 0.024 0.022
RigAnything (Ours) 0.765 0.786 0.765 0.033 0.034 0.019
4.3.2. Connectivity Prediction
Table 3. Connectivity prediction performance comparison on the RigNet dataset. Our method significantly outperforms RigNet across all metrics.
Class. Acc. $\uparrow$ CD-B2B $\uparrow$ ED $\downarrow$
RigNet 0.837 0.001 2.4
RigAnything (Ours) 0.962 0.001 2.3
4.3.3. Skinning Prediction
For skinning prediction performance, we provide a qualitative comparison of our method with RigNet and Blender’s built-in automatic skinning weight calculation, which assigns weights based on the shortest Euclidean distance between mesh vertices and bones in the armature. For a fair comparison, the ground truth skeleton is provided during skinning weight inference. As shown in Fig. 9, our method produces more accurate and consistent skinning weights. In challenging cases where two areas are close in Euclidean space but have a large geodesic distance, our method successfully differentiates these parts and generates consistent skinning weights, whereas the baselines fail.
4.4. Ablation Study
4.4.1. Joint Diffusion
Ours w./o. joint diffusion 0.308 0.277 0.364 0.068 0.059 0.046
Ours w./o. injecting normal 0.559 0.603 0.547 0.053 0.048 0.034
Ours w./o. pose aug. 0.741 0.768 0.732 0.037 0.037 0.022
Ours full model 0.765 0.786 0.765 0.033 0.034 0.019
4.4.2. Normal Injecting
To evaluate the impact of incorporating point normals into the shape tokens, we conducted a comparison experiment without point normals as input. The numerical results in Tab. 4 show a significant decline in skeleton performance when normal information is excluded, highlighting the importance of point normals as geometric information for improving performance.
4.4.3. Online Pose Augmentation
5. Limitation and Future Work
6. Conclusion
Appendix A Final Training Objective
$\mathcal{L}=\mathcal{L}_{\text{joint}}+\mathcal{L}_{\text{connect}}+\mathcal{L}_{\text{skinning}}.$
Appendix B Skeleton Ambiguity
We elaborate on the two types of skeleton ambiguity introduced in the main paper, i.e., the sibling ambiguity and the topology ambiguity. During the BFS ordering process used to traverse and sequentialize the skeleton tree, the ordering of nodes at the same depth in the BFS tree remains undefined. This can lead to the sibling ambiguity, as illustrated in Fig. 10. For example, when the preceding skeleton tokens are 1, 2, and 3, the next joint could either be 4 or 5, each with an equal probability of occurrence. Additionally, an object can have multiple valid skeleton topologies, as shown in Fig. 11, and the model must be capable of capturing these multiple valid configurations. Our method naturally addresses these by modeling the distribution of the next joint based on preceding predictions, which offers a distinct advantage over deterministic approaches in capturing the inherent uncertainty in joint positions.
Figure 11. Two skeleton examples with different skeleton topologies generated on the same shape input.
Appendix C Dataset Details
Our dataset consists of a total of 12,040 shapes, combining 2,354 samples from the RigNet dataset and 9,686 shapes from the Objaverse dataset. From this collection, we utilize 11,217 samples for training and reserve 823 samples for validation and testing. The dataset covers a wide range of object categories and poses, ensuring diversity and robustness in downstream tasks.
C.1. Data Filtering
We apply rigorous data filtering processes to ensure the quality and validity of the dataset. The filtering criteria differ slightly between the RigNet and Objaverse datasets due to their varying quality and size. For the RigNet dataset, which is already high-quality, we apply simple filtering based on the following principles:
  • Shapes with more than 64 joints in their rigging are excluded.
  • Shape are excluded if their skeleton is invalid (e.g. skeleton hierarchy does not form a proper tree structure).
For the Objaverse dataset, which is significantly larger, we employ both manual labeling and automated filtering scripts to ensure consistency and quality. The filtering is performed based on the following principles:
  • Shapes with more than 64 joints in their rigging are excluded.
@@ C.1. Data Filtering
  • Shapes are excluded if their skeleton is invalid (e.g. skeleton hierarchy does not form a proper tree structure).
@@
  • Shapes are excluded if the skeleton does not properly align with the corresponding geometry.
@@
  • Overly simplified or indistinguishable shapes (i.e. shapes consisting of only a few vertices and faces) are excluded.
@@
The Objaverse dataset originally includes 21,622 shapes with rigging annotations. During filtering, we excluded 811 shapes with overly complicated skeletons containing more than 64 joints, which were usually associated with facial or hair rigs. Additionally, 10,471 shapes were filtered out due to low-quality rigging or skinning annotations, leaving us with a refined set of shapes with reliable rigging information.
Figure 12. Distribution of joint numbers across shapes in our dataset.
C.2. Category statistics
@@
For the Objaverse dataset, we also collect the Objaverse category labels and classify the shapes into six categories: humanoid/bipedal, quadruped, insectoid, avian, marine, and other. The ”other” category typically includes manipulable articulated rigid objects found in the dataset (e.g. suitcases, cabinets, etc). Below, we provide the statistics for the number of shapes in each category in Tab. 5:
Table 5. Category statistics of the filtered Objaverse dataset.

### v2

@@ Abstract.
We present RigAnything, a novel autoregressive transformer-based model, which makes 3D assets rig-ready by probabilistically generating joints and skeleton topologies and assigning skinning weights in a template-free manner. Unlike most existing auto-rigging methods, which rely on predefined skeleton templates and are limited to specific categories like humanoid, RigAnything approaches the rigging problem in an autoregressive manner, iteratively predicting the next joint based on the global input shape and the previous prediction. While autoregressive models are typically used to generate sequential data, RigAnything extends its application to effectively learn and represent skeletons, which are inherently tree structures. To achieve this, we organize the joints in a breadth-first search (BFS) order, enabling the skeleton to be defined as a sequence of 3D locations and the parent index. Furthermore, our model improves the accuracy of position prediction by leveraging diffusion modeling, ensuring precise and consistent placement of joints within the hierarchy. This formulation allows the autoregressive model to efficiently capture both spatial and hierarchical relationships within the skeleton. Trained end-to-end on both RigNet and Objaverse datasets, RigAnything demonstrates state-of-the-art performance across diverse object types, including humanoids, quadrupeds, marine creatures, insects, and many more, surpassing prior methods in quality, robustness, generalizability, and efficiency. It achieves significantly faster performance than existing auto-rigging methods, completing rigging in under a few seconds per shape. Please check our website for more details: https://www.liuisabella.com/RigAnything.
@@
submissionid: 181copyright: acmlicensedjournal: TOGjournalyear: 2025journalvolume: 44journalnumber: 4publicationmonth: 8doi: 10.1145/3731149ccs: Computing methodologies Animationccs: Computing methodologies Neural networks
@@ 1. Introduction
Figure 2. Skeleton generation given real images, showing our method generalizes well to real data. More real results in Fig. 10.
@@
Table 1. Feature comparison with other Auto-Rigging tools.
Specifically, we represent the tree-structured skeleton as a sequence by ordering the joints in a breadth-first search (BFS) order, where each joint is defined by a 3D position and a parent index. This autoregressive formulation is particularly suited for skeleton prediction, as it addresses the inherent ambiguity in joint configurations by representing them as a probabilistic distribution. Additionally, by sequentially generating joints and connections without relying on a predefined template, the model supports arbitrary skeleton structures and varying numbers of joints, enabling broad generalization across diverse object categories. Furthermore, while transformer-based autoregressive models are traditionally designed to handle discrete values (Waswani et al., 2017; Brown et al., 2020; Radford et al., 2019), inspired by recent work utilizing autoregressive models for image generation (Li et al., 2024), we adopt a diffusion sampling process to predict the continuously valued joint positions, resulting in superior accuracy. Given the predicted skeleton, we infer the skinning weights by a pair-wise computation. We employ transformer blocks throughout the model to comprehensively capture the global shape structure, as well as the interdependence among all joints and their associated surface points.
@@ 2. Related Work
Figure 3. A single step in our method: The input shape and the previously predicted skeleton sequence are tokenized using two separate tokenizers. These tokens are processed through a chain of autoregressive transformer blocks with a hybrid attention mask. Shape tokens perform self-attention to capture global geometric information, while skeleton tokens attend to all shape tokens and use causal attention within themselves to maintain the autoregressive generation process. After the transformer blocks, a skinning module decodes shape tokens into skinning weights, a joint diffusion module samples the next joint position, and a connectivity module predicts the next joint’s connection to its preceding joints.
@@ 2.1. Automatic Rigging
Rigging is a fundamental technique for animation in computer graphics. Traditional automatic rigging methods, such as Pinocchio (Baran and Popović, 2007), rely on predefined skeletons and optimize their variations to fit a range of characters, with skinning weights determined by analyzing vertex-bone deformation relationships. However, the optimization process is computationally expensive and diminishes the generalizability. Recent advances in deep learning have improved the quality and adaptability of rigging. TARig (Ma and Zhang, 2023) utilizes a template with adaptive joints and a boneflow field to generate skeletons and skinning weights for humanoid characters. Li’s et al (2021) leverages a predefined skeleton template for characters to learn rigging and proposes neural blend shape to enhance deformation quality. However, these methods are confined to humanoid characters in standard poses and rely heavily on predefined templates, limiting their robustness and generalization to diverse objects, poses, and skeleton topologies.
@@ 2.2. Autoregressive Models for 3D
Autoregressive models are a powerful class of probabilistic models widely applied across domains such as natural language processing (Brown et al., 2020; Achiam et al., 2023; Radford et al., 2019) and computer vision (Esser et al., 2021; Parmar et al., 2018; Chen et al., 2020; Li et al., 2024). In 3D tasks, autoregressive models have also demonstrated remarkable potential in areas like shape generation (Yan et al., 2022; Mittal et al., 2022; Ibing et al., 2023; Cheng et al., 2022; Qian et al., 2024) and motion generation (Zhang et al., 2023; Rempe et al., 2021; Han et al., 2024). In 3D shape generation, methods mostly focus on designing effective representations for autoregressive modeling. ShapeFormer (Yan et al., 2022) introduces a sparse representation that quantizes non-empty voxel grids in a predefined order. AutoSDF (Mittal et al., 2022) takes a different approach by modeling the entire space and using randomized sampling orders to enable non-sequential modeling. Octree Transformer (Ibing et al., 2023) introduce octree-based hierarchical shape representations with adaptive compression, significantly reducing sequence lengths. Cheng et al. (Cheng et al., 2022) decompose point clouds into semantically aligned sequences. Argus3D (Qian et al., 2024) utilizes discrete representation learning on a latent vector and scales up the model to improve the quality and versatility of 3D generation. Similarly, autoregressive models have advanced 3D motion generation. MotionVAEs (Ling et al., 2020) uses autoregressive conditional variational autoencoders to learn a latent action space for human movement generation and control. T2M-GPT (Zhang et al., 2023) uses motion VQ-VAE and textural descriptions for human motion generation. HuMoR (Rempe et al., 2021) proposes hierarchical latent variables for realistic motion synthesis. AMD (Han et al., 2024) presents an autoregressive model that iteratively generates complex 3D human motions from long textual descriptions. In this paper, we pioneer the application of autoregressive models to the task of automatic rigging, marking a significant advancement in this domain.
@@ 3. Method
Our goal is to transform a given 3D shape into an animatable asset by generating a plausible skeleton and predicting the corresponding skinning weights. These enable the 3D asset to be articulated under joint transformations using Linear Blend Skinning (LBS). In this section, we first describe two types of skeleton ambiguities in Sec. 3.1. We then present our novel autoregressive model for skeleton generation in Sec. 3.2, followed by our approach to skinning weight prediction in Sec. 3.3. Finally, we provide a detailed description of the model architecture and the overall training objective in Sec. 3.4 and Sec. 3.5.
3.1. Skeleton Ambiguity
We identify two types of skeleton ambiguities that commonly occur in existing pipelines. (1) Sibling ambiguity: The ordering of nodes at the same depth in the skeleton tree is undefined, as illustrated in Fig. 4. For instance, if the skeleton is traversed in breadth-first search (BFS) order and the preceding skeleton tokens are 1, 2, and 3, the next joint could be either 4 or 5, each being equally valid.
(2) Topology ambiguity: An object may have multiple valid skeleton topologies, as shown in Fig. 5. This requires the method to model a distribution over multiple plausible configurations. Our method naturally addresses this ambiguity by modeling the distribution of the next joint based on preceding predictions, which offers a distinct advantage over deterministic approaches in capturing the inherent uncertainty in joint positions.
Figure 5. Examples of different valid skeleton topologies for the same shape.
@@ 3.2. Autoregressive Skeleton Prediction
Autoregressive Modeling. The key component of our method is an autoregressive model for the skeleton prediction to address the ambiguity of skeleton structures and eliminates the need for predefined templates. To convert the tree-structured skeleton to a sequence that can be effectively processed by the autoregressive model, we adopt the BFS to serialize the skeleton to a list:
@@
where $j_{k}\in\mathbb{R}^{3}$ and $p_{k}\in\{1,...,K\}$ denote the 3D position and the parent index of the $k$-th joint respectively. As we adopt the BFS order, $p_{k}<k$ and the first element $(j_{1},p_{1})$ always represents the root joint. The order of joints at the same BFS depth level is non-deterministic. To resolve this ambiguity, we randomly sample the order during training and use generative modeling to cover the uncertainty.
@@
Instead of directly modeling in the original joint space, we map all previously predicted joints and their corresponding parents into a higher-dimensional token space to enhance the model’s expressive capacity. This token space effectively represents the evolving state of the skeleton, capturing its structural and hierarchical information as new joints and connections are incrementally added. Similarly, a sequence of shape tokens is extracted to encapsulate the global structure of the input shape, providing consistent contextual information throughout the modeling process. Denoting the skeleton tokens as $T_{1:k-1}\in\mathbb{R}^{(k-1)\times d}$ and the shape token as $H\in\mathbb{R}^{L\times d}$, where $d$ is the dimension of each token. The prediction targets are reformulated as:
@@
[…]
@@
capturing both the evolving skeleton state and the input shape. The model takes the noisy joint $j^{m}$ as input and predicts the added noise $\epsilon$. The training objective is defined as:
@@ 3.4. Autoregressive Transformer Architecture
Our autoregressive modeling is anchored on a transformer-based architecture, which outputs the shape tokens $H\in\mathbb{R}^{L\times d}$ and skeleton tokens $T_{1:k}\in\mathbb{R}^{k\times d}$ ($0<k\leq K$) that serve as conditional inputs for the autoregressive modeling for skeleton prediction (Sec. 3.2) and skinning prediction (Sec. 3.3). The extraction of these tokens involve two steps: first, referred as the “tokenization” step, an initial shape token and skeleton tokens are lifted from the raw input, this step produces a higher dimensional vector that has sufficient capacity in preparation to capture richer information in the further processing steps in the transformer; subsequently, the transformer process the tokens through a series of attention blocks with carefully crafted attention masking to obtain the final shape and skeleton tokens, which are finally used as inputs to the skeleton and skinning prediction modules in Sec. 3.2 and Sec. 3.3.
@@
3.5. Final Training Objective
(13) $\mathcal{L}=\mathcal{L}_{\text{joint}}+\mathcal{L}_{\text{connect}}+\mathcal{L}_{\text{skinning}}.$
@@ 4.1. Implementation Details
The implemented transformer consists of 12 layers with a hidden dimension of 1024. Following the implementation in (Zhang et al., 2025), each transformer block incorporates a multi-head self-attention layer with 16 heads and a two-layered MLP with a hidden dimension of 4096 and a GeLU activation. We employ Pre-Layer Normalization, Layer Normalization (LN), and residual connections consistent with the reference implementation. During training, we employ a hybrid attention masking strategy: shape tokens perform self-attention to effectively capture geometric information, while skeleton tokens use causal attention, attending only to their ancestor skeleton tokens within the sequence to facilitate auto-regressive generation. Additionally, skeleton tokens attend to all shape tokens. During inference, the network processes shape tokens as input and generates skeleton tokens in an auto-regressive manner. We apply KV caching in the transformer blocks to boost the inference speed.
The joint diffusion process follows (Nichol and Dhariwal, 2021; Li et al., 2024), which has a cosine noise scheduler with 1000 training steps and 50 resampling steps during inference. The denoising MLP is conditioned on the transformer-outputted joint tokens, where these tokens are incorporated into the noise scheduler’s time embedding through AdaLN (Peebles and Xie, 2023) within the Layer Normalization layers.
The fusing module is a two-layer MLP with an input size of 3072 and hidden dimensions of 2048 and 1024. During inference, after obtaining the next joint position via diffusion sampling, a shape tokenizer generates a latent shape token (dimension 1024), which is concatenated with previous context tokens (dimension 1024) and positional embeddings. The fusing module’s output serves as the updated context for connectivity and skinning prediction. Both the connectivity and skinning modules share a similar architecture with the fusing module, except that their input size is 2048.
In the autoregressive skeleton generation process, we maintain a learnable start token (BOS) at the beginning of each skeleton sequence to indicate the start of generation. As we traverse the skeleton in BFS order, the joint sampled from the start token is considered the root joint. The stop condition is determined by checking the parent of the current joint—if a joint’s parent is itself, it indicates that the sequence has reached its end.
@@ 4.2. Dataset
Overview. We utilize both the RigNet dataset (Xu et al., 2020) and the Objaverse dataset (Deitke et al., 2023). The RigNet dataset contains 2,354 high-quality 3D models with ground-truth skeleton and skinning annotations. The Objaverse dataset offers a large collection of 3D models with varying rigging quality. To ensure data reliability, we filtered out 9,686 models with consistent and accurate skeleton and skinning information. Our dataset spans a diverse range of categories, including bipedal, quadrupedal, avian, marine, insectoid, and manipulable rigid objects. For each model, we sample point clouds and face normals from the mesh surface.
@@
During training, we perform online pose augmentation to the input data by randomly deforming the input point clouds using the ground-truth skeleton and skinning. Specifically, we apply random perturbations to each joint, with the maximum rotation angle constrained to 45 degrees, and then deform the original point cloud using the perturbed skeleton and the ground truth skinning. As shown in our ablation study in Sec. 4.4 and Fig. 8, this augmentation strengthens our method’s ability to generalize to objects in different poses.
Data Filtering We apply a thorough data filtering process to ensure the quality and validity of the dataset. Both manual checks and automated scripts are used to maintain consistency and quality. The filtering is based on the following rules:
  • Shapes with more than 64 joints are excluded.
@@
  • Shapes with invalid skeletons (e.g., skeleton hierarchy not forming a proper tree) are excluded.
@@
  • Shapes whose skeletons do not align well with the geometry are excluded.
@@
  • Overly simplified or indistinct shapes (e.g., consisting of very few vertices and faces) are excluded.
@@
The Objaverse dataset originally contains 21,622 shapes with rigging annotations. We exclude 811 shapes with overly complex skeletons (more than 64 joints), which are often related to facial or hair rigs. Additionally, 10,471 shapes are removed due to low-quality rigging or skinning annotations. After filtering, we obtain a refined set of 12,040 shapes with reliable rigging information.
@@
Data Statistics We organize the shapes in the Objaverse dataset into six categories based on their labels: humanoid/bipedal, quadruped, insectoid, avian, marine, and other. The ”other” category mainly includes manipulable articulated rigid objects, such as suitcases, cabinets, and similar items. The number of shapes in each category is summarized in Tab. 2. We also analyze the distribution of joint counts per shape and present the results in Fig. 9. Our dataset contains more shapes with joint counts falling in the ranges $[25,30]$, $[50,55]$, and $[60,64]$.